OpenStudy (anonymous):

Find an equation for the nth term of the arithmetic sequence. -15, -6, 3, 12, ... an = -15 + 9(n + 1) an = -15 x 9(n - 1) an = -15 + 9(n + 2) an = -15 + 9(n - 1)

OpenStudy (raden):

an = a1 + (n-1)d given a1 = -15 and d = 12-3 = 9

OpenStudy (anonymous):

an = -15 + 9(n - 1) ???

OpenStudy (raden):

obviously, you are right :)

OpenStudy (anonymous):

thanks :)

OpenStudy (raden):

welcome :)